OpenStudy (anonymous):

help me solve the system y = x2 - 5 -3x + y = -7

OpenStudy (anonymous):

2x-5=-7+3x > x=2 . y=-1

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

-3x + x^2 - 5 = -7 -3x + x^2 + 2 = 0 x^2 - 3x + 2 = 0 ( x - 2 ) ( x -1 ) = 0 this is what i got

OpenStudy (anonymous):

but u dont need to do ^2 you can solve it without that..

OpenStudy (anonymous):

you need to take the y=.... and put him in the sec equ and then solve it

OpenStudy (anonymous):

and them u get the x and after that u take the x and put him in one of the two equ and that it u find the y.